Two women and five men can together finish a piece of embroidery in four days,while 3 women and 6 men can finish it in 3 days.Find the time taken by 1 woman to finish the work.

Grade:12

1 Answers

Sunil Kumar FP
10 years ago
2W +5M requires 4 days
3W + 6M requires 3 days
apply 4(2W+5M) =3(3W + 6M)
2M=1W

6W complete in 3 days
1Women complete in 6*3=18 days

thanks and regards
sunil kr
askIITian faculty